OpenStudy (anonymous):

find all numbers for which the rational expression is undefined y²-y-5 -------- y²-13y+40

OpenStudy (dumbcow):

solve y^2 -13y +40 =0 factor -8*-5 =40 -8 + -5 = -13 =>(y-8)(y-5) = 0 y = 5 and 8 expression is undefined at these points